OpenStudy (anonymous):

Help!! medal and fan! Kate earned $1800 in 8 years on an investment at a 5% annual simple interest rate. How much did Kate invest?

OpenStudy (campbell_st):

the simple interest formula is \[I = \frac{Prt}{100}\] so substitute what you know \[1800 = \frac{P \times 5 \times 8}{100}\] now just solve for P